OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 3y = –17 2x + 3y = –17 5x + 2y = –4 (3, –1) (–2, 2) (–1, 3) (2, –7)

OpenStudy (anonymous):

you wrote the first equation twice.\[2x + 3y = –17~~~~~~~->~~~~~~~~\times~~ 2~~~~~~~~~~~~->~~~~~~~~~~~~4x+6y=-34\]\[[5x + 2y = –4~~~~~~~->~~~~~~~\times (-3)~~~~~~~~~~->~~~~~~~~~-15x-6y=12\]

OpenStudy (anonymous):

\[add~~~them,~~~you~~~get,~~~~~-11x=-22,~~~~~~~~~x=2\]

OpenStudy (anonymous):

once you get the x, it's obvious which of the choices is correct, but you can still solve for y by plugging 2 for x into one of your equations.

OpenStudy (anonymous):

\[2(2)+3y=-17\]\[4+3y=-17~~~~~~~~~now,~~~subtract~~~4~~~from~~~both~~~~sides.\]\[4+3y-4=-17-4\]\[3y=-21~~~~~~~~~~~~~~~~~~~~y=-7\]

OpenStudy (anonymous):

knowing that according to the first equation if x=2 then y=-7 I'll plug in the values of y and x into the second equation to check. \[5(2)+2(-7)=-4~~~~~~~~~~->~~~~~~~~~~10+(-14)=-4~~~~~~~-4=-4\]